OpenStudy (anonymous):

Working on polynomials and can't seem to get it right. Hope someone can assist me. -4x^3+7x^2-3x-2. x=-2

OpenStudy (anonymous):

do you have to subitue -2 for the xs

OpenStudy (anonymous):

yes I do.

OpenStudy (anonymous):

i got it

OpenStudy (anonymous):

-4x^3+7x^2-3x-2. x=-2 -4(-2)^3+7(-2)^2-3(-2)= 8^3+14^2+6 14^2+8^3+6 is the answer

OpenStudy (anonymous):

try to understand waht i have done i am good at polynimals

OpenStudy (anonymous):

Okay but that is not what the are looking for. It says to evaluate and are looking an answer.

OpenStudy (anonymous):

They also want the answer simplified. I get 3 chances to answer it and then I have to try a different problem. I already used to guesses.

OpenStudy (chrissy):

0

OpenStudy (anonymous):

the answer is 714

OpenStudy (anonymous):

I will try it and tell you what it is. brb

OpenStudy (anonymous):

what

OpenStudy (anonymous):

is 714 your aanswer

OpenStudy (anonymous):

They said the answer was 64

OpenStudy (anonymous):

how

OpenStudy (anonymous):

who

OpenStudy (anonymous):

I am doing my problems online. I work the problem and submit my answer. I get 3 tries then I have to do a different problem.

OpenStudy (anonymous):

After the 3rd try it tells me what the answer is

OpenStudy (anonymous):

try subimating 714

OpenStudy (anonymous):

plz become my fan

OpenStudy (anonymous):

That is what I did. It was my 3rd try so it told me that the answer was 64 and I do not know how they are getting it.